What mass of sodium glycolate should be added to 400 ml of 1.00 M glycolic avid yo produce a buffer solution witj a ph of 4


1
Expert's answer
2021-02-01T04:03:50-0500

for acid buffer mixture:

"\u0440\u041d = \u0440\u041a\u0430\u2013 lg(\u0421acids\/\u0421salts)"

"4 = 3.83\u2013 lg(\u0421acids\/\u0421salts)"

"lg(\u0421acids\/\u0421salts)=-0.17"

"(\u0421acids\/\u0421salts)=0.676"

"(1\/\u0421salts)=0.676"

Сsalts=1.479

"m salts = Сsalts\times V(s)\times М = 1.479\times 0,4\times98.0332=57.996"